# SemShareKV: Efficient KVCache Sharing for Semantically Similar Prompts via Token-Level LSH Matching

### v1

@@ 1 Introduction
Large Language Models (LLMs) have exhibited a strong capability to understand and process human languages, and have been shown to perform comparably to humans in several fields, such as math inference, text memorization, information extraction, story telling (Naveed et al., 2023). Recently released LLMs have significantly advanced in processing and comprehending extremely long prompts. However, this introduces a notable challenge: increased computational demand due to the quadratic time complexity of their Decoder-Only Transformer architecture when handling lengthy text sequences. The issue is further compounded during inference, as the auto-regressive decoding process repeats the computation for each newly generated token (Luohe et al., 2024).
Existing KVCache optimization approaches primarily focus on single-prompt compression through various techniques: Yang et al. (2024a) leverage decaying key-value importance across layers for selective extraction (though with limited small-batch gains), Gim et al. (2024) employ restrictive markup schemas for text chunk reuse, and Yao et al. (2024) propose deviation-based recomputation that requires impractical per-chunk precomputation for long inputs. Crucially, these methods operate within the constrained paradigm of single-prompt optimization, failing to exploit the substantial efficiency potential of cross-prompt cache reuse, a significant oversight given the prevalence of semantically similar queries in real-world applications where shared computational savings could be substantial.
Motivated by this challenge, we aim to address the following research question: Can we reuse the precomputed KVCache for another semantically similar prompt?
@@ 2 Related Work
Prior research on KVCache optimization can be categorized into three key directions: (i) Conventional KVCache Compression, which focuses on reducing the storage and computational overhead of KVCache by applying quantization, pruning, or other compression techniques; (ii) KVCache Sharing, which explores methods to reuse KVCache across different queries or tasks to improve efficiency while maintaining response quality; and (iii) KVCache Reusing, which investigates strategies to adapt and repurpose precomputed KVCache for semantically similar inputs, minimizing redundant computation while preserving model accuracy.
@@ 2.1 Conventional KVCache Compression
To address long-context processing, many works propose optimizing inference by retaining only informative tokens. Token-level compression often uses attention-based token selection (Zhang et al., 2023; Xiao et al., 2024; Li et al., 2024; Yang et al., 2024a; Zhong et al., 2024), low-rank decomposition (Sun et al., 2024), or quantization (Zhang et al., 2024; Wang et al., 2024). Model-level approaches redesign architectures to improve reuse (Sun et al., 2025; Yan et al., 2024), while system-level methods focus on memory and scheduling (Kwon et al., 2023; Sheng et al., 2023). Recent work has highlighted the use of value vectors to facilitate compression (Guo et al., 2024).
@@ 2.2 KVCache Sharing
Some also emphasize reusing portions of the cache for future or similar queries and prompts. For example, PromptCache (Gim et al., 2024) stores text segments that appear frequently on an inference server using a schema, although this approach hampers usability, as users must conform their natural language to the schema format. Mooncake (Qin et al., 2024), KVSharer (Yang et al., 2024b) and MiniCache (Liu et al., ) exploit the high similarity of attention scores among adjacent transformer layers to improve KVCache reuse. By consolidating or sharing Key-Value pairs between similar layers, these methods improve memory efficiency and streamline token processing. However, their approaches are restricted to sharing in the layer or text segment within adjacent layers or the same LLM, limiting the broader applicability; GPTCache (Regmi and Pun, 2024), (Rasool et al., 2024) and (Bang, 2023) utilize similarity search among queries to reuse KVCache. However, they have a high probability of missing a hit and require the entire query to be similar, offering limited flexibility.
@@ 2.3 KVCache Reusing
Limited attention has been directed toward the sharing of KVCache in LLMs. DroidSpeak(Liu et al., 2024b) improves context sharing between fine-tuned LLMs by identifying critical KVCache layers and selectively recomputing them for efficient reuse while maintaining accuracy. LMCache(Cheng et al., 2024) introduces a Knowledge Delivery Network (KDN) to optimize KVCache storage and transfer, allowing cost-effective knowledge injection in LLM inference. CacheBlend(Yao et al., 2024), KVShare(Yang et al., 2025), and EPIC(Hu et al., 2024) rely on exact context matching, which is unsuitable for real user scenarios. SentenceKV suffers from inter-sentence information loss, as noted in the CacheBlend paper. In contrast, SemShareKV introduces RoPE in token matching to address this issue.
@@ 3 Observations and Insights
We present three key insights derived from our experiments on three LLMs: Mistral-7B (Jiang et al., 2023), LLaMA-3.1-8B (Grattafiori et al., 2024), and MPT-7B (Team, 2023). These insights show consistent patterns across different LLMs, supporting the generality of our observations.
@@
When reusing KV caches from semantically similar prompts, we ensure the reused cache maintains high fidelity with fully recomputed caches to prevent performance degradation. To compare the similarity between two KV matrices, we used our augmented MultiNews dataset, where each sample consists of a pair of semantically similar prompts: the Target Prompt, which serves as the primary input to the model, and the Reference Prompt, which acts as the semantically similar counterpart. For each of the aforementioned LLMs, we first computed the KV caches for the prompt pairs independently. Subsequently, we calculated the deviations between the KV caches of the target and reference prompts using the previously mentioned L2 norm. Tokens with the highest 40% deviation were identified as High Deviation (HD) tokens. To further quantify this observation, we compute the Spearman correlation of HD tokens between adjacent layers. As shown in Figure 2, adjacent layers exhibit relatively high consistency in HD token positions.
@@
$S_{\text{total}}=\sum_{i=1}^{n}T_{i}\quad\quad\frac{\sum_{i=k}^{n}T_{i}}{S_{\text{total}}}>\text{Thres}$ (1)
@@
To reduce memory overhead from the KV cache, a key optimization strategy is to remove tokens containing redundant information. Such tokens contribute minimally to the prediction of next tokens during decoding but occupy substantial GPU memory. However, selective token retention risks information loss, necessitating careful trade-offs between memory savings and generation quality. We evaluate three KVCache retention strategies using perplexity: Constant, with equal retention across layers; Exponential Growth, with higher retention in shallow layers; and Exponential Decay, with more retained in deeper layers (Figure C2).
We applied these three retention patterns to LLMs, and utilized perplexity to benchmark the generation performance, shown in Figure C2. For all three LLMs, the Exponential Decay pattern achieves the lowest perplexity score, indicating the best generation performance. This finding further validates that this pattern aligns with how LLMs interpret knowledge from prompts.
@@ 4.1 Relevant Concepts
  • Embedding Cache (E): Contextualized embeddings t capturing fundamental semantic and syntactic relationships (Mikolov et al., 2013), providing the foundational token representations before transformer processing.
@@ 4.2 Model Overview
Once a reference prompt is retrieved, SemShareKV first applies RoPE to the E caches of both the target and reference prompts. Then it uses Locality-Sensitive Hashing (LSH) to match each token in the target prompt to its most similar tokens in the reference prompt. Based on these LSH mappings, the precomputed KVCache of the reference promptrompt is rearranged token by token and injected into LLM transformer layers. On the first transformer layer, all tokens undergo full recomputation. The recomputed outputs are compared with the rearranged cache values via L2 norm, identifying high-deviation tokens for prioritized recomputation in subsequent layers. Simultaneously, the system evicts tokens with the lowest attention scores from recent computations, optimizing KVCache memory usage dynamically.
@@ 4.3 KVCache Sharing Challenge
The primary challenge in cross-prompt KVCache sharing stems from length disparity between prompts. Inspired by (Liu et al., 2024b), we incorporates positional encoding within the E Cache to enable accurate token alignment while preserving contextual relationships.
Specifically, we use LSH to identify, for each token in the target prompt, the most similar token in the reference prompt based on their vector representations. We use Locality-Sensitive Hashing (LSH) for efficient token similarity search. Additional details on LSH are provided in Appendix A.2. This process allowed us to reorder the KVCache of the reference prompt to align with the token sequence of the target prompt. Consequently, the reordered KVCache matches the target prompt’s length, with its key-value pairs entirely derived from the original KVCache of the reference prompt. The LLM uses the reordered KVCache to the target prompt.
@@ 4.3.1 Use Relative Position Encoding to Facilitate Fuzzy Token Match
To address this, we introduce positional encoding into the E cache to enhance fuzzy token matching. Two widely used positional encoding strategies are absolute positional encoding (Vaswani et al., 2017), which embeds explicit position information, and relative positional encoding (Su et al., 2024), which captures positional relationships between tokens. In our work, we incorporate Rotary Position Embedding (RoPE) into the E cache and evaluate its impact. Specifically, RoPE is applied to the non-contextual embeddings (E cache) of both the reference and target prompts. Then, LSH is used to match each token in the target prompt’s E cache to the most similar token in the reference prompt’s E cache. This step is crucial because RoPE introduces position-sensitive information, allowing the same token at different positions to carry distinct semantics, enabling LSH to achieve more accurate token-level matching.
@@
Beyond the initial tokens, a key difference emerges: without positional encoding, many tokens in the target prompt map to the initial tokens, whereas with positional encoding, they align more accurately with later tokens. We interpret this as a manifestation of "attention sink", a phenomenon in self-attention mechanisms where a significant portion of attention scores is consistently assigned to initial tokens, regardless of their actual relevance to the task (Xiao et al., 2023; Fei et al., 2025). Incorporating positional encoding into the E cache effectively mitigates this issue, leading to more accurate token matching and improved performance.
@@ 4.3.2 Impact from Position Encoding in KVCache
The second challenge is that LSH-based token rearrangement disrupts position encoding in the precomputed KV cache, affecting the KV matrices from the prefill phase. Previous studies (Gao et al., 2024) have discussed the impact of different position encoding strategies, suggesting that RoPE can be excluded from KVCache by applying it after storage, as illustrated in Figure 5. In evaluating the role of RoPE in the E cache, we compared three configurations of KVCache: (i) with position encoding, the standard setting where RoPE is applied before storing KVCache; (ii) without position encoding, where RoPE is not applied during storage; and (iii) without position encoding but reapplied, where RoPE is omitted during storage but reapplied after LSH-based reordering. Ideally, the rearranged KVCache should closely match the ground-truth KVCache for the target prompt. To quantify the deviation, we compute the $L_{2}$ norm between the rearranged and ground-truth caches. As shown in Figure 7, KV cache with position encoding has the lowest deviation, followed by the version without position encoding, while the configuration with reapplied RoPE gives the highest deviation. This highlights the importance of storing KV pairs after RoPE is applied. Ensuring consistent position encoding between the E and KV cache is essential for the LLM to fully leverage them and achieve optimal generation quality.
@@ 4.3.3 Recomputation Strategy
We divide the layers into two groups: the first and subsequent layers. Given that LLMs tend to focus more on later tokens (Liu et al., 2024a; Yang et al., 2024a), we categorize tokens into Cold ($c$) and Hot ($h$) using a dynamic ratio $\boldsymbol{r_{dynamic}}$ from Attention Recovery with a threshold of 55%, meaning $\boldsymbol{r_{dynamic}}\%$ of tokens with the highest cumulative attention are selected as Hot, and the rest as Cold. The total number of recomputed tokens is defined as $\mbox{Recomputed}=\omega_{c}\cdot c+\omega_{h}\cdot h$, where $\boldsymbol{\omega_{c}}=0.1$ and $\boldsymbol{\omega_{h}}=0.5$ in the SemShareKV setting.
@@ 5.1 Experiments Setup
We select a diverse set of datasets covering a broad range of tasks. For Q&A, we use WikiHow (Koupaee and Wang, 2018) and Qasper (Dasigi et al., 2021). For summarization, we include MultiNews (Bai et al., 2023) (multi-document), SAMsum (Gliwa et al., 2019) (dialogue), and BookSum, PubMed, and BigPatent (Kwan et al., 2023), which represent narrative, scientific, and patent documents, respectively; all three are single-document summarization. For code completion, we use LLC (Guo et al., 2023). For multiple-choice Q&A, we evaluate on MMLU (Hendrycks et al., 2021b, a).
We compared SemShareKV against three baselines: (i) Fully Recompute: standard inference using the unmodified model from the Transformers library, where the entire prompt is input without any KVCache reuse; (ii) SnapKV (Li et al., 2024): a KVCache management method that accelerates the prefill phase by efficient caching but does not compress the KVCache; (iii) H2O (Zhang et al., 2023): a dynamic KVCache eviction strategy that compresses KV memory by prioritizing important tokens, but does not optimize the prefill phase. ChatGPT said: We use a modified H2O compressing 10% of the cache per layer, with SnapKV and H2O as baselines for prefill optimization and KVCache compression. Experiments ran on a single A100 GPU with standard attention. Implementation details are in Appendix D.

### v2

@@ 1 Introduction
Large Language Models (LLMs) have exhibited a strong capability to understand and process human languages, and have been shown to perform comparably to humans in several fields, such as math inference, text memorization, information extraction, story telling (naveed2023comprehensive). Recently released LLMs have significantly advanced in processing and comprehending extremely long prompts. However, this introduces a notable challenge: increased computational demand due to the quadratic time complexity of their Decoder-Only Transformer architecture when handling lengthy text sequences. The issue is further compounded during inference, as the auto-regressive decoding process repeats the computation for each newly generated token (luohe2024keep).
Existing KV cache optimization approaches primarily focus on single-prompt compression through various techniques: (yang2024pyramidinfer) leverage decaying KV importance across layers for selective extraction (though with limited small-batch gains), (gim2024prompt) employ restrictive markup schemas for text chunk reuse, and (yao2024cacheblend) propose deviation-based recomputation that requires impractical per-chunk precomputation for long inputs. Crucially, these methods operate within the constrained paradigm of single-prompt optimization, failing to exploit the substantial efficiency potential of cross-prompt cache reuse, a significant oversight given the prevalence of semantically similar queries in real-world applications where shared computational savings could be substantial.
Motivated by this challenge, we aim to address the following research question: Can we reuse the precomputed KV cache for prompts that are semantically similar?
@@ 2 Related Work
Prior research on KV cache optimization can be categorized into three key directions: (i) Conventional KVCache Compression, which focuses on reducing the storage and computational overhead of KV cache by applying quantization, pruning, or other compression techniques; (ii) KVCache Sharing, which explores methods to reuse KV cache across different queries or tasks to improve efficiency while maintaining response quality; and (iii) KVCache Reusing, which investigates strategies to adapt and re-purpose precomputed KV cache for semantically similar inputs, minimizing redundant computation while preserving model accuracy.
@@ 2.1 Conventional KVCache Compression
To address long-context processing, many works propose optimizing inference by retaining only informative tokens. Token-level compression often uses attention-based token selection (zhang2023h2o; xiao2024duoattention; li2024snapkv; yang2024pyramidinfer; zhong2024zigzagkv), low-rank decomposition (sun2024shadowkv), or quantization (zhang2024pqcache; wang2024squeezeattention). Model-level approaches redesign architectures to improve reuse (sun2025you; yan2024recurformer), while system-level methods focus on memory and scheduling (kwon2023efficient; sheng2023flexgen). Recent work has highlighted the use of value vectors to facilitate compression (guo2024attention).
@@ 2.2 KVCache Sharing
Some also emphasize reusing portions of the cache for future or similar queries and prompts. For example, PromptCache (gim2024prompt) stores text segments that appear frequently on an inference server using a schema, although this approach hampers usability, as users must conform their natural language to the schema format. Mooncake (qin2024mooncake), KVSharer (yang2024kvsharer) and MiniCache (liu2405minicache) exploit the high similarity of attention scores among adjacent transformer layers to improve KV cache reuse. By consolidating or sharing KV pairs between similar layers, these methods improve memory efficiency and streamline token processing. However, their approaches are restricted to sharing in the layer or text segment within adjacent layers or the same LLM, limiting the broader applicability; GPTCache (regmi2024gpt), (rasool2024llms) and (bang2023gptcache) utilize similarity search among queries to reuse KV cache. However, they have a high probability of missing a hit and require the entire query to be similar, offering limited flexibility.
@@ 2.3 KVCache Reusing
Limited attention has been directed toward the sharing of KV cache in LLMs. DroidSpeak (liu2024droidspeak) improves context sharing between fine-tuned LLMs by identifying critical KV cache layers and selectively recomputing them for efficient reuse while maintaining accuracy. LMCache (cheng2024large) introduces a Knowledge Delivery Network (KDN) to optimize KV cache storage and transfer, allowing cost-effective knowledge injection in LLM inference. CacheBlend (yao2024cacheblend), KVShare (yang2025kvshare), and EPIC (hu2024epic) rely on exact context matching, which is unsuitable for real user scenarios. SentenceKV (Zhu2025SentenceKVEL) suffers from inter-sentence information loss, as noted in the CacheBlend. In contrast, SemShareKV introduces RoPE in token matching to address this issue.
@@ 3 Observations and Insights
We present three key insights derived from our experiments on three LLMs: Mistral-7B (jiang2023mistral), LLaMA-3.1-8B (grattafiori2024llama), and MPT-7B (MosaicML2023Introducing). These insights show consistent patterns across different LLMs, supporting the generality of our observations.
@@
When reusing KV caches from semantically similar prompts, we ensure the reused cache maintains high fidelity with fully recomputed caches to prevent performance degradation. To compare the similarity between two KV matrices, we used our augmented MultiNews (bai2023longbench) dataset, where each sample consists of a pair of semantically similar prompts: the Target Prompt, which serves as the primary input to the model, and the Reference Prompt (Ref Prompt), which acts as the semantically similar counterpart. For each of the aforementioned LLMs, we first computed the KV caches for the prompt pairs independently. Subsequently, we calculated the deviations between the KV caches of the target and reference prompts using the previously mentioned $L_{2}$ norm. Tokens with the highest 40% deviation were identified as High Deviation (HD) tokens. To further quantify this observation, we computed the Spearman correlation of HD tokens between adjacent layers. As shown in Figure 2, adjacent layers exhibit relatively high consistency in HD token positions.
@@
[…]
@@
To reduce memory overhead from the KV cache, a key optimization strategy is to remove tokens containing redundant information. Such tokens contribute minimally to the prediction of next tokens during decoding but occupy substantial GPU memory. However, selective token retention risks information loss, necessitating careful trade-offs between memory savings and generation quality. We evaluate three KV cache retention strategies using perplexity: Constant, with equal retention across layers; Exponential Growth, with higher retention in shallow layers; and Exponential Decay, with more retained in deeper layers. More details of the three retention patterns can be found in Figure C2 in the Appendix.
We applied these three retention patterns to LLMs and evaluated generation performance using perplexity. Across all three models, the Exponential Decay pattern achieves the lowest perplexity, indicating the best performance and suggesting this pattern aligns with how LLMs interpret prompt knowledge. Notably, MPT-7B exhibits a spike in AR at the final layer due to its structural differences from Mistral and Llama3.2: unlike the RoPE-based models, MPT-7B employs ALiBi positional encoding and a pre-norm design. ALiBi biases intermediate layers toward recent tokens, while the final layer compensates by attending to more distant tokens, resulting in broader attention patterns and higher AR values.
@@ 4.1 Relevant Concepts
  • Embedding Cache (E): Contextualized embeddings capturing fundamental semantic and syntactic relationships (Mikolov2013EfficientEO), providing the fundamental token representations before transformer processing.
@@ 4.2 Model Overview
Once a reference prompt is retrieved, SemShareKV first applies RoPE to the E caches of both the target and reference prompts. Then it uses LSH to match each token in the target prompt to its most similar tokens in the reference prompt. Based on these LSH mappings, the precomputed KV cache of the reference prompt is rearranged token by token and injected into LLM transformer layers. On the first transformer layer, all tokens undergo full recomputation. The recomputed outputs are compared with the rearranged cache values via $L_{2}$ norm, identifying high-deviation tokens for prioritized recomputation in subsequent layers. Simultaneously, the system evicts tokens with the lowest attention scores from recent computations, optimizing KV cache memory usage dynamically.
@@ 4.3 KVCache Sharing Challenge
The primary challenge in cross-prompt KV cache sharing stems from length disparity between prompts. Inspired by (liu2024droidspeak), we incorporates positional encoding within the E Cache to enable accurate token alignment while preserving contextual relationships.
Specifically, we use LSH to identify, for each token in the target prompt, the most similar token in the reference prompt based on their vector representations. We use LSH for efficient token similarity search. This process allowed us to reorder the KV cache of the reference prompt to align with the token sequence of the target prompt. Consequently, the reordered KV cache matches the target prompt’s length, with its key-value pairs entirely derived from the original KV cache of the reference prompt. The LLM uses the reordered KV cache to the target prompt. Additional details on LSH are provided in Appendix A.2.
@@ 4.3.1 Use Relative Position Encoding to Facilitate Fuzzy Token Match
To address this, we introduce positional encoding into the E cache to enhance fuzzy token matching. Two widely used positional encoding strategies are absolute positional encoding (Vaswani2017AttentionIA), which embeds explicit position information, and relative positional encoding (su2024roformer), which captures positional relationships between tokens. In our work, we incorporate Rotary Position Embedding (RoPE) into the E cache and evaluate its impact. Specifically, RoPE is applied to the non-contextual embeddings (E cache) of both the reference and target prompts. Then, LSH is used to match each token in the target prompt’s E cache with the most similar token in the reference prompt’s E cache. This step is crucial because RoPE introduces position-sensitive information, allowing the same token at different positions to carry distinct semantics, enabling LSH to achieve more accurate token-level matching.
@@
Beyond the initial tokens, a key difference emerges: without positional encoding, many tokens in the target prompt map to the initial tokens, whereas with positional encoding, they align more accurately with later tokens. We interpret this as a manifestation of "attention sink", a phenomenon in self-attention mechanisms where a significant portion of attention scores is consistently assigned to initial tokens, regardless of their actual relevance to the task (xiao2023efficient; fei2025efficient). Incorporating positional encoding into the E cache effectively mitigates this issue, leading to more accurate token matching and improved performance.
@@ 4.3.2 Impact from Position Encoding in KVCache
The second challenge is that LSH-based token rearrangement disrupts position encoding in the precomputed KV cache, affecting the KV matrices from the prefill phase. Previous studies (Gao2024CostEfficientLL) have discussed the impact of different position encoding strategies, suggesting that RoPE can be excluded from KV cache by applying it after storage, as illustrated in Figure 5. In evaluating the role of RoPE in the E cache, we compared three configurations of KV cache: (i) With position encoding, the standard setting where RoPE is applied before storing KV cache; (ii) Without position encoding, where RoPE is not applied during storage; and (iii) Wthout position encoding but reapplied, where RoPE is omitted during storage but reapplied after LSH-based reordering. Ideally, the rearranged KV cache should closely match the ground-truth KV cache for the target prompt. To quantify the deviation, we compute the $L_{2}$ norm between the rearranged and ground-truth caches. As shown in Figure 7, KV cache with position encoding has the lowest deviation, followed by the version without position encoding, while the configuration with reapplied RoPE gives the highest deviation. This highlights the importance of storing KV pairs after RoPE is applied. Ensuring consistent position encoding between the E and KV cache is essential for LLMs to fully leverage them and achieve optimal generation quality.
@@ 4.3.3 Recomputation Strategy
We divide the layers into two groups: the first layer and subsequent layers. Given that LLMs tend to focus more on later tokens (liu2024lost; yang2024pyramidinfer), we categorize tokens into Cold ($c$) and Hot ($h$) using a dynamic ratio $\boldsymbol{r_{dynamic}}$ from Attention Recovery with a threshold of 55%, meaning $\boldsymbol{r_{dynamic}}\%$ of tokens with the highest cumulative attention are selected as Hot, and the rest as Cold. The total number of recomputed tokens is defined as $\mbox{Recomputed}=\omega_{c}\cdot c+\omega_{h}\cdot h$, where $\boldsymbol{\omega_{c}}=0.1$ and $\boldsymbol{\omega_{h}}=0.5$ in the SemShareKV.
@@ 5.1 Experiments Setup
We select a diverse set of datasets covering a broad range of tasks. For Q&A, we use WikiHow (koupaee2018wikihow) and Qasper (Dasigi2021ADO). For summarization, we include MultiNews (bai2023longbench) (multi-document), SAMsum (gliwa-etal-2019-samsum) (dialogue), and BookSum, PubMed, and BigPatent (kwan_m4le_2023), which represent narrative, scientific, and patent documents, respectively; all three are single-document summarization. For code completion, we use LLC (guo2023longcoder). For multiple-choice Q&A, we evaluate on MMLU (hendryckstest2021; hendrycks2021ethics).
We compared SemShareKV against four baselines: (i) Fully Recompute: standard inference using the unmodified model from the Transformers library, where the entire prompt is input without any KV cache reuse; (ii) SnapKV (li2024snapkv): a KV cache management method that accelerates the prefill phase by efficient caching but does not compress the KV cache; (iii) PyramidKV (cai2024pyramidkv): PyramidKV leverages the Pyramidal Information Funneling pattern in LLMs by dynamically adjusting KV cache sizes across layers, allocating more in lower layers and less in higher ones; (iv) H2O (zhang2023h2o): a dynamic KV cache eviction strategy that compresses KV memory by prioritizing important tokens, but does not optimize the prefill phase. We use a modified H2O compressing 10% of the cache per layer, with SnapKV and H2O as baselines for prefill optimization and KV cache compression. Our framework is based on a simple LRU mechanism to manage precomputed caches, and SemShareKV can be easily integrated with other cache management strategies. The experiments were run on a single A100 GPU with standard attention. The implementation details are in Appendix D.